# GLHF: General Learned Evolutionary Algorithm Via Hyper Functions

### v1

@@ 4.4 Visualization Analysis
Figure 9 displays $S^{t}$ for an in-depth analysis of the LMM strategy (refer to Appendix Figure 14-19 for additional details). Key observations and conclusions include: 1) Generally, superior individuals receive higher weights during LMM, showcasing GPOM’s ability to balance exploration and exploitation as the population converges. 2) Across diverse function problems, GPOM dynamically generates optimization strategies, highlighting its adaptability and contributing to robust generalization. 3) Disadvantaged individuals exhibit a more uniform weight distribution, potentially aiding in their escape from local optima and enhancing algorithm convergence.
@@
We visually examine the LCM strategy, presenting the results in Fig. 10 (refer to Appendix Figure 20-25 for additional details). LCM displays the capacity to adaptively generate diverse strategies for individuals across different ranks in the population, revealing distinct patterns among tasks and rankings. Notably, top-ranking individuals within the top 20, such as those ranked 1st, 5th, and 18th, exhibit a flexible crossover strategy. The dynamic adjustment of crossover probability with population evolution aids in preserving dominant genes and facilitating escape from local optima. Conversely, lower-ranking individuals show an increasing overall probability of crossover, promoting exploration of disadvantaged individuals and enhancing the algorithm’s exploration capability. LCM proficiently generates adaptive crossover strategies across tasks, individuals, and convergence stages, significantly boosting both convergence and exploration capabilities.
@@ 5 Conclusions
We present GPOM, a novel General Pretrained Optimization Model, addressing inefficiencies in existing methods. GPOM, a population-based pretrained BBO model, leverages Differential Evolution (DE) mechanisms with mutation, crossover, and selection operators. Modules like LMM and LCM equip GPOM with advanced optimization capabilities. Pretraining GPOM ensures robust convergence and generalization. Evaluation on BBOB and robot control tasks demonstrates GPOM’s superiority over other pretrained optimization models, even in high-dimensional scenarios. GPOM’s adaptive optimization strategy and remarkable generalization mark significant advancements, addressing key challenges in existing models.
Further research avenues include: 1) Designing enhanced loss functions to optimize GPOM for both population convergence and diversity, thus enhancing algorithm performance. 2) The correlation between the training set and test problem may impact GPOM’s performance during testing. To mitigate this, employing a "pre-training + fine-tuning" approach could be beneficial. This involves pre-training GPOM on various tasks followed by fine-tuning on the specific target problem with few samples to improve its efficacy in solving the target task.
6 Potential Impact
Appendix A Mutation Strategy in DE
@@ Appendix B Landscape Features of TF1-Tf8
  • TF6: Multimodal, Separable, Asymmetrical, Local optima’s number is huge
  • TF7: Multi-modal, Non-separable, Rotated
@@ C.1 BBOB
BBOB (Finck et al., 2010; Hansen et al., 2021) is a widely researched and recognized collection of benchmark test problems to evaluate the performance of optimization algorithms. The dataset consists of a series of high-dimensional continuous optimization functions, including single-peak, multi-peak, rotated, and distorted functions, as well as some functions with specific properties such as Lipschitz continuity and second-order differentiability.
@@ C.2 Robot Control Tasks
We test the performance of GPOM on two complex robot control tasks.
@@ C.2.1 Bipedal Walker
The continuous control task Bipedal Walker (Brockman et al., 2016), implemented within the Box2D physics engine, has been designed to test the ability of walking agents to navigate varying terrain by controlling their joints and maintaining balance. The challenge requires the agent to learn efficient walking strategies that enable it to traverse the intended path without falling or deviating from its trajectory. The robot’s state comprises a range of variables, including the hull angle speed, angular velocity, horizontal speed, vertical speed, joint positions and angular speeds, legs contact with the ground, and lidar rangefinder measurements. The robot’s actions involve determining motor speed values in the range of [-1, 1] for each of the four joints at the hips and knees. The performance of the agent is evaluated through a reward system, whereby it receives points for moving forward, with a maximum of 300+ points awarded upon successfully reaching the end of the designated course. However, the penalty of -100 points is imposed if the robot loses balance and falls. Furthermore, applying motor torque incurs a small cost in terms of points. The score accrued by the agent serves as a measure of its optimal performance. The Bipedal Walker task represents a challenging and dynamic environment that effectively evaluates the walking and balance control abilities of agents. As such, it provides a valuable benchmark for testing and comparing different reinforcement learning algorithms for robotic locomotion.
@@ C.2.2 Enduro
Enduro (Brockman et al., 2016) is one of the classic reinforcement learning environments provided by OpenAI Gym. It is a driving racing game based on the Atari 2600 game. In this environment, your goal is to drive as far as possible by controlling the car. The Enduro game is set on an endless highway where you need to avoid other vehicles and overtake as many other vehicles as possible within a limited time. You can avoid collisions with other vehicles by moving your car left and right, and be careful to control your speed to avoid accidents. The game rewards you based on how far you drive, so your goal is to learn a good driving strategy to maximize the distance traveled.
@@ Appendix D Baselines
Our core is the population-based pre-training BBO algorithm, so we do not compare with non-population methods such as Bayesian optimization methods. Moreover, Bayesian optimization methods are difficult to deal with continuous optimization problems of more than 100 dimensions. We do not use LLM-based approaches (Romera-Paredes et al., 2023; Meyerson et al., 2023; Liu et al., 2023a; Yang et al., 2023; Lehman et al., 2023; Ma et al., 2023; Chen et al., 2023; Nasir et al., 2023) as baselines because they can only be used for a specific type of task.
Heuristic Population-based BBO Algorithm. DE(DE/rand/1/bin) (Das & Suganthan, 2010), ES(($\mu$,$\lambda$)-ES), L-SHADE (Tanabe & Fukunaga, 2014), and CMA-ES (Hansen, 2016), where DE (Das & Suganthan, 2010) and ES are implemented based on Geatpy (Jazzbin, 2020), CMA-ES and IPOP-CMA-ES are implemented by cmaes [^1] https: //github.com/CyberAgentAILab, and L-SHADE is implemented by pyade222 https://github.com/xKuZz/pyade. The reasons for choosing these baselines are the following:
@@
Pretrained BBO Algorithm. We chose three state-of-the-art meta-learn BBO algorithms for comparison with GPOM.
  • LES (Lange et al., 2023b): A recently proposed learnable ES. It uses a data-driven approach to discover new ES with strong generalization performance and search efficiency.
  • LGA (Lange et al., 2023a): A recently proposed learnable GA that discovers new GA in a data-driven manner. The learned algorithm can be applied to unseen optimization problems, search dimensions, and evaluation budgets.
  • We train GPOM on $TS$. During training, the maximum number of evolution generations is 100, $n=100$ and the problem dimension is set to 10.
@@ Appendix E Parameters and Training Dataset
The primary control parameters of CMA-ES and L-SHADE are automatically adjusted. For LGA and LES, we utilized the optimal parameters provided by the authors without modifications. Other hyperparameters were tuned using grid search to identify the optimal combinations, and multiple experiments were conducted accordingly. Detailed parameter settings are presented in Table 2. Each experiment reports the mean and standard deviation of the results from various sets of experiments, with a consistent population size of 100 across all trials.
@@
GPOM $d_{m}=1000$ Standard Settings for GPOM (M).
@@
DE $\textit{F}=0.5$ we determine this hyperparameter between $[0.1,0.9]$ using a grid search, with a step of 0.1. (Jazzbin, 2020).
@@
LES
Table 3: GPOM parameters of different architectures and architecture settings.
@@
TF6(Rastrigin) $\sum\limits_{i=1}^{d}(z_{i}^{2}-10\cos(2\pi z_{i})+10)$ $x\in[-5,5],\omega\in[-2.5,2.5]$
TF7(Griewank) $\sum\limits_{i=1}^{d}{\frac{z_{i}^{2}}{4000}}-\prod_{i=1}^{d}\cos(\frac{z_{i}}{\sqrt{i}})+1$ $x\in[-600,600],\omega\in[-300,300]$
@@ F.1 BBOB Test
Table 5: BBOB RESULT. GPOM is trained on TF1-Tf4 with $d$ =10. The best results are indicated in bold, and the suboptimal results are underlined.
$d$ F GPOM ES DE CMA-ES LSHADE LES LGA
@@ F.2 BBOB Test With Optimal Solution Disturbed
We further test the performance of the algorithm on the BBOB. Here, the optimal solution of each function is randomly disturbed, that is, $\mathbf{x}^{*}=\mathbf{x}_{opt}+\mathbf{z}$, where $\mathbf{x}^{*}$ represents the optimal solution after disturbing, $\mathbf{x}_{opt}$ represents the original optimal solution, $\mathbf{x}_{opt}$ is a vector obtained by random sampling and $\mathbf{z}\in[-1,1]^{d}$. The results are displayed in Table 6. We found that the performance of GPOM can still dominate other algorithms when the function optimal solution is disturbed.
@@ F GPOM ES DE CMA-ES LSHADE LES LGA
Figure 12: The log convergence curves of GPOM and other baselines. It shows the convergence curve of these algorithms on functions in BBOB with $d=30$.
@@
Figure 13: The log convergence curves of GPOM and other baselines. It shows the convergence curve of these algorithms on the functions in BBOB with $d=100$.
Appendix G Results of Analysis Study
@@ F NO LMM NO LCM NO MASK UNTRAINED GPOM
Table 10: Results of GPOMs of different sizes on BBOB tests ($d=100$). The best results are indicated in bold, and the suboptimal results are underlined.

### v2

@@ 4.4 Visualization Analysis
Figure 8 displays $S^{t}$ for an in-depth analysis of the LMM strategy (refer to Appendix Figure 15-20 for additional details). Key observations and conclusions include: 1) Generally, superior individuals receive higher weights during LMM, showcasing POM’s ability to balance exploration and exploitation as the population converges. 2) Across diverse function problems, POM dynamically generates optimization strategies, highlighting its adaptability and contributing to robust generalization. 3) Disadvantaged individuals exhibit a more uniform weight distribution, potentially aiding in their escape from local optima and enhancing algorithm convergence.
@@
We visually examine the LCM strategy, presenting the results in Fig. 9 (refer to Appendix Figure 21-26 for additional details). LCM displays the capacity to adaptively generate diverse strategies for individuals across different ranks in the population, revealing distinct patterns among tasks and rankings. Notably, top-ranking individuals within the top 20, such as those ranked 1st, 5th, and 18th, exhibit a flexible crossover strategy. The dynamic adjustment of crossover probability with population evolution aids in preserving dominant genes and facilitating escape from local optima. Conversely, lower-ranking individuals show an increasing overall probability of crossover, promoting exploration of disadvantaged individuals and enhancing the algorithm’s exploration capability. LCM proficiently generates adaptive crossover strategies across tasks, individuals, and convergence stages, significantly boosting both convergence and exploration capabilities.
@@ 5 Conclusions
We present POM, a novel Pretrained Optimization Model designed to address the inefficiencies of existing methods in zero-shot optimization. Evaluation on BBOB and robot control tasks demonstrates POM’s superiority over other black-box optimizers, particularly in high-dimensional scenarios. Additionally, POM excels in solving few-shot optimization problems. Future research avenues include designing enhanced loss functions to optimize POM for both population convergence and diversity, thereby improving overall algorithm performance. In addition, the limitations of model scale and time performance deserve further study (see Appendix I for details).
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 62206205 and 62471371, in part by the Young Talent Fund of Association for Science and Technology in Shaanxi, China under Grant 20230129, in part by the Guangdong High-level Innovation Research Institution Project under Grant 2021B0909050008, and in part by the Guangzhou Key Research and Development Program under Grant 202206030003.
Appendix A Preliminaries
A.1 Genetic Algorithms
The crossover, mutation, and selection operators form the basic framework of GAs. GA starts with a randomly generated initial population. Then, genetic operations such as crossover and mutation will be carried out. After the fitness evaluation of all individuals in the population, a selection operation is performed to identify fitter individuals to undergo reproduction to generate offspring. Such an evolutionary process will be repeated until specific predefined stopping criteria are satisfied.
Crossover
The crossover operator generates a new individual $\mathbf{x}^{c}_{i}\in\mathbb{R}^{d}$ by Eq. (17), and $cr$ is the probability of the crossover operator.
$x^{c}_{k}=\left\{\begin{matrix}x^{i}_{k}&rand(0,1)<cr\\ x^{j}_{k}&otherwise\end{matrix}\right.$ (17)
where $k\in[1,\cdots,d]$, $i$ and $j$ $\in[1,2,\dots,n]$ ($i\neq j$). $d$ represents the dimension of the problem. $x_{k}^{i}$ and $x_{k}^{j}$ represent the k-th element of $\mathbf{x}^{i}$ and $\mathbf{x}^{j}$ respectively. This operator is commonly conducted on $n$ individuals; $n$ represents the population size. After an expression expansion, we reformulate Eq. (17) as $\sum_{i=1}^{n}\mathbf{x}_{i}\mathbf{W}_{i}^{c}$ [46]. $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the ith individual in $\mathbf{X}$, where $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{n}\}$ is a population. $\mathbf{W}_{i}^{c}\in\mathbb{R}^{d\times d}$ is the diagonal matrix. If $\mathbf{W}_{i}^{c}$ is full of zeros, the $i$th individual has no contribution.
Mutation
The mutation operator brings about random changes in the population. Specifically, an individual $\mathbf{x}_{i}$ in the population goes through the mutation operator to form the new individual $\mathbf{x}_{i}^{m}$, formulated as follows:
$x_{k}^{m}=\left\{\begin{matrix}rand(l_{k},u_{k})&rand(0,1)<mr\\ x_{k}^{c}&otherwise\end{matrix}\right.$ (18)
where $mr$ is the probability of mutation operator and $k\in[1,\cdots,d]$. $x_{k}^{m}$ and $x_{k}^{c}$ represent the k-th element of $\mathbf{x}^{m}$ and $\mathbf{x}^{c}$ respectively. Similarly, Equation (18) can be reformulated as $\mathbf{x}_{i}^{c}\mathbf{W}_{i}^{m}$, where $\mathbf{W}_{i}^{m}\in\mathbb{R}^{d\times d}$ is the diagonal matrix.
Selection
We introduce the binary tournament mating selection operator in Eq. (19). The selection operator survives individuals of higher quality for the next generation until the number of individuals is chosen. As shown in Eq. (19),
$p_{i}=\left\{\begin{matrix}1&f(\mathbf{{x}}_{i})<f(\mathbf{{x}}_{k})\\ 0&f(\mathbf{{x}}_{i})>f(\mathbf{{x}}_{k})\end{matrix}\right.,\ \ (\mathbf{{x}}_{i},\mathbf{{x}}_{k})\in\mathbf{{X}}\cup\mathbf{X}^{m}$ (19)
where $p_{i}$ reflects the probability that $\mathbf{{x}}_{i}$ is selected for the next generation, and $\mathbf{X}^{m}=\{\mathbf{x}_{1}^{m},\mathbf{x}_{2}^{m},\cdots,\mathbf{x}_{n}^{m}\}$.
A.2 Mutation Strategy in DE
The core components of the optimization model include modules that generate solutions and modules that select solutions. GA and DE basically include crossover modules, mutation modules and selection modules. The evolutionary strategy represented by CMA-ES needs to sample a population from a certain distribution (such as Gaussian distribution), and further select individuals to update this distribution. In this paper, we design parameterized trainable LMM and LCM as modules for generating solutions. The function of LMM is to generate a candidate population, and LCM further performs crossover between the candidate population and the original population to obtain the offspring population.
@@ Appendix B Landscape Features of TF1-TF8
  • TF6: Multi-modal, Non-separable, Rotated
  • TF7: Multimodal, Separable, Asymmetrical, Local optima’s number is huge
@@ C.1 BBOB
BBOB [59, 55] is a widely researched and recognized collection of benchmark test problems to evaluate the performance of optimization algorithms. The dataset consists of a series of high-dimensional continuous optimization functions, including single-peak, multi-peak, rotated, and distorted functions, as well as some functions with specific properties such as Lipschitz continuity and second-order differentiability.
@@ C.2 Robot Control Tasks
We test the performance of POM on two complex robot control tasks.
@@ C.2.1 Bipedal Walker
The continuous control task Bipedal Walker [58], implemented within the Box2D physics engine, has been designed to test the ability of walking agents to navigate varying terrain by controlling their joints and maintaining balance. The challenge requires the agent to learn efficient walking strategies that enable it to traverse the intended path without falling or deviating from its trajectory. The robot’s state comprises a range of variables, including the hull angle speed, angular velocity, horizontal speed, vertical speed, joint positions and angular speeds, legs contact with the ground, and lidar rangefinder measurements. The robot’s actions involve determining motor speed values in the range of [-1, 1] for each of the four joints at the hips and knees. The performance of the agent is evaluated through a reward system, whereby it receives points for moving forward, with a maximum of 300+ points awarded upon successfully reaching the end of the designated course. However, the penalty of -100 points is imposed if the robot loses balance and falls. Furthermore, applying motor torque incurs a small cost in terms of points. The score accrued by the agent serves as a measure of its optimal performance. The Bipedal Walker task represents a challenging and dynamic environment that effectively evaluates the walking and balance control abilities of agents. As such, it provides a valuable benchmark for testing and comparing different reinforcement learning algorithms for robotic locomotion.
@@ C.2.2 Enduro
Enduro [58] is one of the classic reinforcement learning environments provided by OpenAI Gym. It is a driving racing game based on the Atari 2600 game. In this environment, your goal is to drive as far as possible by controlling the car. The Enduro game is set on an endless highway where you need to avoid other vehicles and overtake as many other vehicles as possible within a limited time. You can avoid collisions with other vehicles by moving your car left and right, and be careful to control your speed to avoid accidents. The game rewards you based on how far you drive, so your goal is to learn a good driving strategy to maximize the distance traveled.
@@
In these two tasks we conduct 10 sets of experiments, each set of experiments consists of 5 independent runs. We finally take the best results of each set of experiments to calculate the mean and standard deviation.
@@ Appendix D Baselines
Our core is the population-based pre-training BBO algorithm, so we do not compare with non-population methods such as Bayesian optimization methods. Moreover, Bayesian optimization methods are difficult to deal with continuous optimization problems of more than 100 dimensions. We do not use LLM-based approaches [28, 29, 30, 31, 32, 34, 35, 36] as baselines because they can only be used for a specific type of task.
Heuristic Population-based BBO Algorithm. DE(DE/rand/1/bin) [54], ES(($\mu$,$\lambda$)-ES), L-SHADE [22], and CMA-ES [12], where DE [54] and ES are implemented based on Geatpy [60], CMA-ES and IPOP-CMA-ES are implemented by cmaes [^1] https: //github.com/CyberAgentAILab, and L-SHADE is implemented by pyade222 https://github.com/xKuZz/pyade. The reasons for choosing these baselines are the following:
@@
Pretrained BBO Algorithm. We chose three state-of-the-art meta-learn BBO algorithms for comparison with POM.
  • LES [25]: A recently proposed learnable ES. It uses a data-driven approach to discover new ES with strong generalization performance and search efficiency.
  • LGA [10]: A recently proposed learnable GA that discovers new GA in a data-driven manner. The learned algorithm can be applied to unseen optimization problems, search dimensions, and evaluation budgets.
  • We train POM on $TS$. During training, the maximum number of evolution generations is 100, $n=100$ and the problem dimension is set to 10.
@@ Appendix E Parameters and Training Dataset
The primary control parameters of CMA-ES and L-SHADE are automatically adjusted. For LGA and LES, we utilized the optimal parameters provided by the authors without modifications. Other hyperparameters were tuned using grid search to identify the optimal combinations, and multiple experiments were conducted accordingly. Detailed parameter settings are presented in Table 1. Each experiment reports the mean and standard deviation of the results from various sets of experiments, with a consistent population size of 100 across all trials. All experiments are performed on a device with GeForce RTX 3090 24G GPU, Intel Xeon Gold 6126 CPU and 64G RAM.
@@
POM $d_{m}=1000$ Standard Settings for POM (M).
@@
DE $\textit{F}=0.5$ we determine this hyperparameter between $[0.1,0.9]$ using a grid search, with a step of 0.1. [60].
@@
LES All parameters We use the pre-trained optimal parameters provided by the authors.
Table 2: POM parameters of different architectures and architecture settings.
@@
TF6(Griewank) $\sum\limits_{i=1}^{d}{\frac{z_{i}^{2}}{4000}}-\prod_{i=1}^{d}\cos(\frac{z_{i}}{\sqrt{i}})+1$ $x\in[-600,600],\omega\in[-300,300]$
TF7(Rastrigin) $\sum\limits_{i=1}^{d}(z_{i}^{2}-10\cos(2\pi z_{i})+10)$ $x\in[-5,5],\omega\in[-2.5,2.5]$
@@ F.1 BBOB Test
Table 4: BBOB RESULT. POM is trained on TF1-TF5 with $d$ =10. The best results are indicated in bold, and the suboptimal results are underlined.
$d$ F POM ES DE CMA-ES LSHADE LES LGA
@@ F.2 BBOB Test With Optimal Solution Disturbed
We further test the performance of the algorithm on the BBOB. Here, the optimal solution of each function is randomly disturbed, that is, $\mathbf{x}^{*}=\mathbf{x}_{opt}+\mathbf{z}$, where $\mathbf{x}^{*}$ represents the optimal solution after disturbing, $\mathbf{x}_{opt}$ represents the original optimal solution, $\mathbf{x}_{opt}$ is a vector obtained by random sampling and $\mathbf{z}\in[-1,1]^{d}$. The results are displayed in Table 5 and Figure 10. We found that the performance of POM can still dominate other algorithms when the function optimal solution is disturbed.
@@ F POM ES DE CMA-ES LSHADE LES LGA
Figure 12: The log convergence curves of POM and other baselines. It shows the convergence curve of these algorithms on functions in BBOB with $d=30$.
@@
Figure 13: The log convergence curves of POM and other baselines. It shows the convergence curve of these algorithms on the functions in BBOB with $d=100$.
Appendix G Compare with TurBO
We compared POM and Bayesian optimization algorithms, finding that Bayesian optimization converges very slowly on high-dimensional problems. TurBO [56], noted for its fast convergence and strong performance [57], was used as a benchmark. Although TurBO requires substantial time for 10,000 evaluations, POM completes the same task in under one second. Therefore, we plotted the convergence curves of TurBO and POM with only 3,100 evaluations. As shown in Figure 14, POM demonstrates significant performance advantages over TurBO in most cases.
(a) F1
(b) F2
(c) F3
(d) F4
(e) F5
(f) F6
(g) F7
(h) F8
(i) F9
(j) F10
(k) F11
(l) F12
(m) F13
(n) F14
(o) F15
(p) F16
(q) F17
(r) F18
(s) F19
(t) F20
(u) F21
(v) F22
(w) F23
(x) F24
Figure 14: The log convergence curves of POM and TurBO. It shows the convergence curve of these algorithms on functions in BBOB with $d=100$.
Appendix H Results of Analysis Study
@@ F NO LMM NO LCM NO MASK UNTRAINED POM
Table 9: Results of POMs of different sizes on BBOB tests ($d=100$). The best results are indicated in bold, and the suboptimal results are underlined.